OpenStudy (anonymous):

-4(x+3)=16 how would this be worked out?

OpenStudy (anonymous):

Okay, what is the first step?

OpenStudy (anonymous):

Distributive property multiply -4 to x and 3 in the parentheses

OpenStudy (anonymous):

Okay, so now what does the equation look like after doing that?

OpenStudy (anonymous):

-12x=16

OpenStudy (anonymous):

Not quite. Let start with -4 times x. What do you get?

OpenStudy (anonymous):

-4x right?

OpenStudy (anonymous):

Yep, okay, now -4 times 3.

OpenStudy (anonymous):

-12

OpenStudy (anonymous):

Yep, so now we have -4x - 12 = 16 What's next?

OpenStudy (anonymous):

gotta get x by itself so -4x on both sides

OpenStudy (anonymous):

no plus -4x not minus

OpenStudy (anonymous):

on both sides

OpenStudy (anonymous):

Let's start by getting rid of the -12 that way we don't have to move the x.

OpenStudy (anonymous):

right that makes more sense lol

OpenStudy (anonymous):

-4x=32

OpenStudy (anonymous):

then u divide by -4 on both sides and u get x=-8

OpenStudy (anonymous):

Careful with the addition: -4x - 12 + 12 = 16 + 12 -4x = ?

OpenStudy (anonymous):

the 12+ 12 cancels out 12+16=32 -4x=32 divide -4 on each side -4/-4 cancles out 32/-4=-8 right??

OpenStudy (anonymous):

12 + 16 = 28

OpenStudy (anonymous):

But you're right on with the dividing by -4.

OpenStudy (anonymous):

opps got in a hurry leme fix it

OpenStudy (anonymous):

No worries, I do it all the time. (Probably thought 16 and 16 instead, in that case you'd be right with the 32.)

OpenStudy (anonymous):

lol x=-7

OpenStudy (anonymous):

Yep! You got it!